OpenStudy (anonymous):

Calculate the concentration (in molarity) of NaOH solution if 25.0mL of the solution are needed to neutralize 17.4 mL of a 0.132M HCL solution

OpenStudy (rogue):

If you want to neutralize a certain amount of HCl, you need to have the same amount of NaOH. You can use the titration equation to find the molarity of NaOH required.\[M_A V_A = M_B V_B\]The A is for acid and B is for base by the way.\[17.4 mL \times 0.132 M = 25.0 mL \times M_B\]\[M_B = \frac {17.2 mL \times 0.132 M}{25.0 mL} = 0.0908 M\]So you need 0.091 M of NaOH to neutralize that amount of HCl.